OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t2 + 640t. After how many seconds does the projectile take to reach its maximum height? Show your work for full credit.

OpenStudy (mathmate):

No calculus: h(t)=-16(t^2-40t) =-16(t-20)^2+6400 So maximum (vertex) occurs at t=20. With calculus: h'(t)=-32t+640=0 Solve for t=20

OpenStudy (anonymous):

Awesome! Thank you :)

OpenStudy (mathmate):

You're welcome! :)